Two charges q1 and q2 are placed such that, the electric field at one-fourth of the way from
q1 to q2 is zero. Determine the ratio q1:q2
(1) 1:9 (2) 2:1 (3) 1:16 (4) 16:1
1
Expert's answer
2019-04-17T12:13:46-0400

At this point we have


"E_1=E_2"

"\frac{kq_1}{(\frac{d}{4})^2}=\frac{kq_2}{(\frac{3d}{4})^2}"

"9q_1=q_2"

Answer: (1) 1:9